Let A and B be two events in a sample space S such that P(A) = 0.6, P(B) = 0.4, and P(A ∩ B) = 0.18. Find the probabilities below.
(a) P(A|B)
(b) P(B|A)

Answer by stanbon(75887)

Let A and B be two events in a sample space S such that P(A) = 0.6, P(B) = 0.4, and P(A ∩ B) = 0.18. Find the probabilities below.
(a) P(A|B) = P(A and B)/P(B) = 0.18/0.4 = 0.45
---------------------------------------------------
(b) P(B|A) = P(A and B)/P(A) = 0.18/0.6 = 0.3
==================================================
cheers,
Stan H.